how can you turn this peoblem into a linear equation? 4x=2y+8
Answer by ewatrrr(24785)

 
Hi,
4x=2y+8 , 4x-2y = 8 is the standard form
Solving for y gives one the slope-intercept form, which is useful:
y = 2x - 4, y = mx + b, slope m = 2 and y-intercept is -4.